OpenStudy (anonymous):

What is the standard form equation of the line passing through (3, -2) with a slope of 3 over 5?

OpenStudy (anonymous):

a) 5x - 3y = 19 b) 5x - 3y = -19 c) 3x - 5y = 19 d) 3x - 5y = -19

OpenStudy (anonymous):

m=3/5 (3,-2) Y=mx+c ...sub into this equation

OpenStudy (anonymous):

-2= 3/5(3)+c

OpenStudy (anonymous):

like this?

OpenStudy (anonymous):

yep then find c once you have found c...multiply by 5 throughout

OpenStudy (anonymous):

idk how to do it... i got the equation but idk how to solve

OpenStudy (anonymous):

y=3/5 X-19/5 5Y= 3/5 (5) X- 19/5 (5)

OpenStudy (anonymous):

i dont know how to multiply with the fraction.

OpenStudy (shamim):

formula needed for it\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

@danny562 3/5 X 5/1 ...like that u cancel the 5

OpenStudy (shamim):

given\[x _{1}=3,y _{1}=-2,m=\frac{ 3 }{ 5 }\]

OpenStudy (shamim):

so the equation of line will b\[y-(-2)=\frac{ 3 }{ 5 }(x-3)\]

OpenStudy (anonymous):

so the answer is c

OpenStudy (shamim):

ya

OpenStudy (anonymous):

3x-5y=19

OpenStudy (shamim):

ya

OpenStudy (anonymous):

@danny562 yes....yay!!